# Taxonomic classification with maximal exact matches in KATKA kernels and minimizer digests

### 10 Feb 2024 v1

@@ 4 Experiments
For each genus , we simulated 500 reads of 200 base pairs each by choosing a random starting location in the reference sequence for and mutating 1% percent of the bases uniformly across the read to simulate sequencing error. For each read and each index, we found all the read’s longest MEMs and checked whether all their [first, last] ranges contained only the ID of the reference sequence for . Figure 10 shows the index sizes and true-positive rates over all 500 000 simulated reads. Clearly, we can achieve significant compression while only slightly decreasing the true-positive rate, especially with KATKA kernels of minimizer digests: for example, with and our index took 56.5 MiB and achieved a true-positive rate of 74.3%, compared to 287.9 MiB and 78.6% with an index for the full dataset, better than the tradeoffs we achieve with kernelization or minimizers alone.
@@ 5 Conclusions and future work
Figure 10 strongly confirms our conjecture from Subsection 2.4 that KATKA kernels of minimizer digests can inherit the strengths of both. In the near future we plan experiment also with varying the width of minimizers (for simplicity, in this paper we always used 3-mers) and to measure also the speedups we can achieve. (Searching over minimizer digests is usually significantly faster than searching over original texts, both because some characters are not represented in the digests and because we use a backward step for each minimizer rather than for each character, incurring fewer cache misses.) Later, we plan to incorporate indexing KATKA kernels of minimizer digests to build MEM tables — with more sophisticated classifications that take advantage of all the information in those tables — into a full pipeline for taxonomic classification of reads.

### 04 Apr 2024 v2

@@ 4 Experiments
For each genus , we simulated 500 reads of 200 base pairs each by choosing a random starting location in the reference sequence for and mutating 1% percent of the bases uniformly across the read to simulate sequencing error. For each read and each index, we found all the read’s longest MEMs and checked whether all their [first, last] ranges contained only the ID of the reference sequence for . Figure 10 shows the index sizes and true-positive rates over all 500 000 simulated reads, and Figure 11 shows an average search time per read.
Clearly, we can achieve significant compression while only slightly decreasing the true-positive rate and without paying a penalty in search time, especially with KATKA kernels of minimizer digests: for example, with and our index took 56.5 MiB and achieved a true-positive rate of 74.3%, compared to 287.9 MiB and 78.6% with an index for the full dataset; our index is also slightly faster than the index for the full dataset. This is a better compression/accuracy/speed tradeoff than we achieve with either kernelization or minimizers alone.
If we are willing to sacrifice the true-positive rate moderately, we can increase and also achieve significant speedups. FM-indexes over minimizer digests are known to be usually significantly faster than indexes over the original datasets, both because some characters are not represented in the digests and because we use a backward step for each minimizer rather than for each character, incurring fewer cache misses. Interestingly, however, we achieved slightly better speedups with both kernelization and minimizers than with either separately.
@@ 5 Conclusions and future work
Figure 10 strongly confirms our conjecture from Subsection 2.4 that KATKA kernels of minimizer digests can inherit the strengths of both. In the near future we plan to experiment also with varying the width of minimizers (for simplicity, in this paper we always used 3-mers). Later, we plan to incorporate indexing KATKA kernels of minimizer digests to build MEM tables — with more sophisticated classifications that take advantage of all the information in those tables — into a full pipeline for taxonomic classification of reads.
@@
Acknowledgments
Many thanks to Sana Kashgouli and Finlay Maguire for helpful discussions. D.D. funded by CTU grant No. SGS23/205/OHK3/3T/18 and by ROBOPROX reg. no. CZ.02.01.01/00/22_008/0004590. O.Y.A. and B.L. funded by NIH grants R35GM139602 & R01HG011392 and NSF grant DBI-2029552. T.G. funded by NSERC Discovery Grant RGPIN-07185-2020. G.M. funded by the Italian Ministry of Health, POS 2014-2020, project ID T4-AN-07, CUP I53C22001300001, by INdAM-GNCS Project CUP E53C23001670001 and by PNRR ECS00000017 Tuscany Health Ecosystem, Spoke 6 CUP I53C22000780001. G.N. funded by Basal Funds FB0001, ANID, Chile.